# DiCoFlex: Model-agnostic diverse counterfactuals with flexible control

### v1

@@ 4.2 Evaluating multiple explanations
Lending DiCoFlex 1.000 0.999 0.668 0.412 0.840 0.476 13.118
Club s-DiCoFlex 1.000 0.997 0.704 0.635 0.737 0.453 11.373
@@
Adult DiCoFlex 1.000 0.998 0.581 0.515 0.498 2.156 2.036
s-DiCoFlex 1.000 1.000 0.373 0.568 0.350 1.812 0.379
@@
Bank DiCoFlex 1.000 0.971 0.717 0.499 0.786 0.276 3.808
s-DiCoFlex 1.000 0.894 0.584 0.480 0.673 0.247 3.006
@@
Default DiCoFlex 1.000 0.986 0.459 0.419 0.831 0.038 62.427
s-DiCoFlex 1.000 0.988 0.489 0.514 0.787 0.034 61.593
@@
The comparison with baselines for diverse counterfactual explanations (DiCE, CCHVAE) presented in Table 1 confirms that DiCoFlex returns more diverse explanations than competitive methods, achieving superior hypervolume scores. In terms of plausibility evaluation, both versions of DiCoFlex consistently outperform other baseline methods as indicated by the LOF measure (except the GMC dataset). Furthermore, the results reveal that the proposed method is capable of producing counterfactual explanations with a low proximity score on continuous attributes.
However, impressive diversity, plausibility, and proximity come at the price of slightly worse sparsity. Although s-DiCoFlex obtained the lowest $\epsilon$-sparsity on continuous features in three out of five cases, the categorical sparsity is higher than the one returned by DiCE and CCHAVE. It is apparent that these two methods switch categorical variables less often than continuous ones, which is usually caused by the applied optimization scheme and can be seen as their drawback. Observe that the use of s-DiCoFlex allows us to balance the level of sparsity between categorical and continuous features.
@@
4.3 DiCoFlex vs. s-DiCoFlex
As mentioned, lowering the value of $p$ in the $L_{p}$ norm for DiCoFlex directly results in lowering the sparsity measure for continuous attributes. As a side effect, it allows for larger changes in values of categorical variables to produce valid counterfactuals. An interesting consequence is that s-DiCoFlex generates more plausible samples than DiCoFlex (see the LOF score), which may be attributed to the way we model the distribution $q$ used for scoring counterfactual candidates; see Section 3.3. For low $p=0.01$, we restrict the support of $q$ to a small region of data space, where only a small number of coordinates are likely to be modified. Increasing this value leads to neighbors located in a larger and more diverse region since we allow for modifying all coordinates without extra penalty. Consequently, DiCoFlex with $p=2$ can sample out-of-distribution counterfactuals more often than s-DiCoFlex with $p=0.01$. In practice, dynamic change of $p$ in inference allows users to examine multiple counterfactual candidates and select the one that suits their expectations best.
@@ 4.4 Comparison with single counterfactual methods
To present a broader view on the performance of DiCoFlex, we include an additional comparison with commonly used methods that generate a single explanation (ReViSE, TABCF, Wachter) (see Table 1). We note that DiCoFlex and s-DiCoFlex still demonstrate impressive performance in terms of LOF score in this comparison. They both achieve the lowest LOF on Adult and Bank datasets and are comparable to the best-performing ReViSE method on the Default dataset. Moreover, they obtained the best continuous proximity measure in 3 out of 5 cases. This means that even though DiCoFlex produces multiple diverse counterfactuals for each instance, most of them satisfy the highest requirements used in counterfactual analysis.
@@ 4.5 Runtime of counterfactuals generation
A key advantage of our approach is that once trained, DiCoFlex can be used to generate multiple counterfactual explanations in a single forward pass. To illustrate this gain, we measure the time every method needs to generate a single counterfactual for 100 data points. Figure 2 confirms that both variants of DiCoFlex achieve generation times of $0.12\pm 0.05$ seconds, outperforming other baseline methods by orders of magnitude. Optimization-based methods demonstrate substantially higher computational costs, with DiCE, ReViSE, and TABCF all requiring over 1000 seconds on average. This efficiency enables real-time exploration of diverse counterfactual explanations with dynamic constraint adjustment, addressing a critical limitation of existing approaches.
@@ 4.6 Actionability Control
DiCoFlex is capable of imposing additional constraints when generating counterfactuals without retraining the model. As demonstrated in Section 4.3 and Table 1, the use of the $L_{p}$ norm with small $p$ allows to minimize the modifications of the continuous features and improve plausibility. In this experiment, we show the effect of introducing actionability constraints, which prevents severe modification of user-defined features.
@@
For illustration, we train DiCoFlex on the Adult dataset with four mask configurations, which cover one or two features, as shown in Table 2. For instance, the use of "mask 1" should block changes of "Capital Gain" and "Capital Loss" attributes. As can be observed in the result, the application of every individual mask indeed decreases the sparsity level compared to the unconstrained version when no mask is used. In consequence, this flexible control enables users to dynamically adjust both sparsity and actionability, generating counterfactuals satisfying different domain-specific requirements without requiring model retraining. Appendix G contains additional evaluation of these models.
@@ 5 Conclusion
We presented DiCoFlex, a model-agnostic conditional generative framework for counterfactual explanations that addresses key limitations in existing approaches. By leveraging normalizing flows, our method enables counterfactual generation without constant access to the classification model, provides inference-time control over constraints, and efficiently produces multiple diverse explanations in a single forward pass. Empirical evaluations on five benchmark datasets demonstrate that DiCoFlex outperforms the state-of-the-art methods in diversity and validity while achieving superior plausibility and computational efficiency. Although optimization-based methods may achieve better sparsity in some cases, our approach offers a more balanced trade-off between competing explanation criteria.
Appendix A Dataset Descriptions

### v2

@@ 4.2 Evaluating multiple explanations
Lending DiCoFlex (p=2.0) 1.000 0.999 0.668 0.412 0.840 0.476 13.118
Club DiCoFlex (p=0.01) 1.000 0.997 0.704 0.635 0.737 0.453 11.373
@@
Adult DiCoFlex (p=2.0) 1.000 0.998 0.581 0.515 0.498 2.156 2.036
DiCoFlex (p=0.01) 1.000 1.000 0.373 0.568 0.350 1.812 0.379
@@
Bank DiCoFlex (p=2.0) 1.000 0.971 0.717 0.499 0.786 0.276 3.808
DiCoFlex (p=0.01) 1.000 0.894 0.584 0.480 0.673 0.247 3.006
@@
Default DiCoFlex (p=2.0) 1.000 0.986 0.459 0.419 0.831 0.038 62.427
DiCoFlex (p=0.01) 1.000 0.988 0.489 0.514 0.787 0.034 61.593
@@
The comparison with baselines for diverse counterfactual explanations (DiCE, CCHVAE) presented in Table˜1 confirms that DiCoFlex returns more diverse explanations than competitive methods, achieving superior hypervolume scores. In terms of plausibility evaluation, both versions of DiCoFlex consistently outperform other baseline methods as indicated by the LOF measure (except the GMC dataset). Furthermore, the results reveal that the proposed method is capable of producing counterfactual explanations with a low proximity score on continuous attributes.
However, impressive diversity, plausibility, and proximity come at the price of slightly worse sparsity. Although DiCoFlex (p=0.01) obtained the lowest $\epsilon$-sparsity on continuous features in three out of five cases, the categorical sparsity is higher than the one returned by DiCE and CCHAVE. It is apparent that these two methods switch categorical variables less often than continuous ones, which is usually caused by the applied optimization scheme and can be seen as their drawback. Observe that the use of DiCoFlex (p=0.01) allows us to balance the level of sparsity between categorical and continuous features.
@@
4.3 DiCoFlex (p=2.0) vs. DiCoFlex (p=0.01)
As mentioned, lowering the value of $p$ in the $L_{p}$ norm for DiCoFlex directly results in lowering the sparsity measure for continuous attributes. As a side effect, it allows for larger changes in values of categorical variables to produce valid counterfactuals. An interesting consequence is that DiCoFlex (p=0.01) generates more plausible samples than DiCoFlex (p=2.0) (see the LOF score), which may be attributed to the way we model the distribution $q$ used for scoring counterfactual candidates; see Section˜3.3. For low $p=0.01$, we restrict the support of $q$ to a small region of data space, where only a small number of coordinates are likely to be modified. Increasing this value leads to neighbors located in a larger and more diverse region since we allow for modifying all coordinates without extra penalty. Consequently, DiCoFlex with $p=2$ can sample out-of-distribution counterfactuals more often than DiCoFlex (p=0.01) with $p=0.01$. In practice, dynamic change of $p$ in inference allows users to examine multiple counterfactual candidates and select the one that suits their expectations best.
@@ 4.4 Comparison with single counterfactual methods
To present a broader view on the performance of DiCoFlex, we include an additional comparison with commonly used methods that generate a single explanation (ReViSE, TABCF, Wachter) (see Table˜1). We note that DiCoFlex and DiCoFlex (p=0.01) still demonstrate impressive performance in terms of LOF score in this comparison. They both achieve the lowest LOF on Adult and Bank datasets and are comparable to the best-performing ReViSE method on the Default dataset. Moreover, they obtained the best continuous proximity measure in 3 out of 5 cases. This means that even though DiCoFlex produces multiple diverse counterfactuals for each instance, most of them satisfy the highest requirements used in counterfactual analysis.
@@ 4.5 Runtime of counterfactuals generation
A key advantage of our approach is that once trained, DiCoFlex can be used to generate multiple counterfactual explanations in a single forward pass. To illustrate this gain, we measure the time every method needs to generate a single counterfactual for 100 data points. Figure˜2 confirms that both variants of DiCoFlex achieve generation times of $0.12\pm 0.05$ seconds, outperforming other baseline methods by orders of magnitude. Optimization-based methods demonstrate substantially higher computational costs, with DiCE, ReViSE, and TABCF all requiring over 1000 seconds on average. This efficiency enables real-time exploration of diverse counterfactual explanations with dynamic constraint adjustment, addressing a critical limitation of existing approaches.
@@ 4.6 Actionability Control
DiCoFlex is capable of imposing additional constraints when generating counterfactuals without retraining the model. As demonstrated in Section˜4.3 and Table˜1, the use of the $L_{p}$ norm with small $p$ allows to minimize the modifications of the continuous features and improve plausibility. In this experiment, we show the effect of introducing actionability constraints, which prevents severe modification of user-defined features.
@@
For illustration, we train DiCoFlex on the Adult dataset with four mask configurations, which cover one or two features, as shown in Table˜2. For instance, the use of "mask 1" should block changes of "Capital Gain" and "Capital Loss" attributes. As can be observed in the result, the application of every individual mask indeed decreases the sparsity level compared to the unconstrained version when no mask is used. In consequence, this flexible control enables users to dynamically adjust both sparsity and actionability, generating counterfactuals satisfying different domain-specific requirements without requiring model retraining. Appendix˜I contains additional evaluation of these models.
4.7 Sensitivity Analysis of Sparsity Parameter $p$
We further analyze the sensitivity of DiCoFlex to the sparsity control parameter $p$ that governs the $L_{p}$ norm used during neighbor selection (Section 3.4). This parameter determines the strength of sparsity constraints applied to continuous features and thus influences the trade-off between proximity, sparsity, and diversity.
Table 3 reports the performance of DiCoFlex across several $p$ values on the Adult dataset. Lower values (e.g., $p=0.01$) favor sparse modifications, yielding counterfactuals that stay closer to the original instances, while larger values (e.g., $p=1$ or $2$) permit more flexible changes that improve diversity at the cost of proximity.
Table 3: Effect of sparsity parameter $p$ on DiCoFlex performance on the Adult dataset. Lower proximity and LOF, higher validity and hypervolume indicate better performance.
Classif. Proximity Sparsity $\epsilon$-sparsity LOF Hypervol.
Model prob. ↑ cont. ↓ cat. ↓ cont. ↓ log ↓ log ↑
DiCoFlex (p=0.01) 1.000 0.373 0.568 0.441 1.9119 0.9023
DiCoFlex (p=0.08) 0.996 0.501 0.553 0.454 1.9680 1.0801
DiCoFlex (p=0.25) 0.995 0.551 0.541 0.501 2.2380 1.3780
DiCoFlex (p=1.0) 0.996 0.577 0.537 0.589 2.6349 1.9491
DiCoFlex (p=2.0) 0.998 0.581 0.515 0.601 2.6614 1.8862
Across all $p$ values, DiCoFlex maintains perfect validity, confirming robustness of the learned conditional flow. As $p$ increases, proximity gradually worsens while diversity (hypervolume) improves, reflecting a smooth proximity–sparsity trade-off. The consistent trend in LOF indicates stable plausibility across regimes.
These findings demonstrate that a single DiCoFlex model, trained with conditioning on $p$, can smoothly adapt its behavior to user preferences at inference time—ranging from highly sparse to diverse counterfactuals—without retraining or model modification.
@@ 5 Conclusion
We presented DiCoFlex, a model-agnostic conditional generative framework for counterfactual explanations that addresses key limitations in existing approaches. By leveraging normalizing flows, our method enables counterfactual generation without constant access to the classification model, provides inference-time control over constraints, and efficiently produces multiple diverse explanations in a single forward pass. Our approach is grounded in rigorous theoretical foundations: we prove formal guarantees for validity, proximity bounds, in-distribution plausibility, and diversity preservation (Appendix A). Empirical evaluations on five benchmark datasets demonstrate that DiCoFlex outperforms the state-of-the-art methods in diversity and validity while achieving superior plausibility and computational efficiency. Although optimization-based methods may achieve better sparsity in some cases, our approach offers a more balanced trade-off between competing explanation criteria.
Acknowledgments and Disclosure of Funding
Oleksii Furman and Maciej Zieba’s work was supported by the National Science Centre (Poland) Grant No. 2024/55/B/ST6/02100. The research of U. Movsum-zada and P. Marszałek was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169. The work of M. Śmieja was supported by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. Some experiments were performed on servers purchased with funds from the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Krakow.
Appendix A Theoretical Considerations
This section provides theoretical analysis of the diversity guarantees provided by DiCoFlex. We establish a formal lower bound on the expected diversity of counterfactuals generated by our learned conditional normalizing flow.
A.1 Preliminary Lemma
Lemma A.1 (Upper bound on difference of expected values).
Let $p$ and $q$ be two distributions over the same finite set $\mathcal{X}$, and let the total variation distance be defined as $\|p-q\|_{\text{TV}}=\frac{1}{2}\sum_{x\in\mathcal{X}}|p(x)-q(x)|$. For any function $f:\mathcal{X}\to\mathbb{R}$, we have:
$\left|\sum_{x\in\mathcal{X}}f(x)(p(x)-q(x))\right|\leq 2\cdot\max_{x\in\mathcal{X}}|f(x)|\cdot\|p-q\|_{\text{TV}}.$ (9)
Proof.
By the triangle inequality and the definition of total variation distance:
$\left|\sum_{x\in\mathcal{X}}f(x)(p(x)-q(x))\right|$ $\leq\sum_{x\in\mathcal{X}}|f(x)|\cdot|p(x)-q(x)|$ (10)
$\leq\max_{x\in\mathcal{X}}|f(x)|\cdot\sum_{x\in\mathcal{X}}|p(x)-q(x)|$ (11)
$=2\cdot\max_{x\in\mathcal{X}}|f(x)|\cdot\|p-q\|_{\text{TV}}.$ (12)
A.2 Main Theoretical Result
Theorem A.2 (Diversity Preservation in DiCoFlex).
Let $p_{\theta}(x^{\prime}|x,y^{\prime},p,m)$ be the learned conditional normalizing flow and $\hat{q}(x^{\prime}|x,y^{\prime},d_{p,m})$ be the empirical distribution defined in Equation (6). Under assumptions (A1)-(A3) stated below, for $n$ samples $\{x^{\prime}_{i}\}_{i=1}^{n}$ drawn from $p_{\theta}$, the expected diversity satisfies:
$\mathbb{E}_{p_{\theta}}[D(\{x^{\prime}_{i}\}_{i=1}^{n})]\geq\sigma_{\min}(1-K^{1-n})-\Delta\sqrt{2\epsilon},$ (13)
where $D(\{x^{\prime}_{i}\}_{i=1}^{n})=\min_{i\neq j}\|x^{\prime}_{i}-x^{\prime}_{j}\|$ denotes the minimum pairwise distance among the generated counterfactuals.
Assumptions
  • (A1)
    Minimum neighborhood separation: The neighborhood $N(x,y^{\prime},d_{p,m},K)$ contains $K$ distinct points with minimum pairwise distance $\sigma_{\min}>0$. Specifically, for any $x_{i},x_{j}\in N(x,y^{\prime},d_{p,m},K)$ with $i\neq j$,
    $\|x_{i}-x_{j}\|\geq\sigma_{\min}.$ (14)
  • (A2)
    Training convergence: The normalizing flow is trained sufficiently close to the empirical distribution, satisfying:
    $D_{\text{KL}}(\hat{q}(x^{\prime}|x,y^{\prime},d_{p,m})\|p_{\theta}(x^{\prime}|x,y^{\prime},p,m))\leq\epsilon$ (15)
    for all $(x,y^{\prime},p,m)$ in the training distribution.
  • (A3)
    Bounded support: The support of $p_{\theta}(\cdot\mid x,y^{\prime},p,m)$ lies in a set of finite diameter $\Delta$, i.e., for any $x^{\prime},x^{\prime\prime}$ in the support,
    $\|x^{\prime}-x^{\prime\prime}\|\leq\Delta.$ (16)
    This implies that $0\leq D(\{x^{\prime}_{i}\}_{i=1}^{n})=\min_{i\neq j}\|x^{\prime}_{i}-x^{\prime}_{j}\|\leq\Delta$.
Proof.
Step 1: Lower bound on empirical diversity. Consider the empirical distribution $\hat{q}(x^{\prime}|x,y^{\prime},d_{p,m})$ which samples uniformly from $K$ distinct neighbors. When sampling $n$ points with replacement from these $K$ points, the probability of selecting at least two distinct points is:
$P_{n}(K)=1-K\cdot\left(\frac{1}{K}\right)^{n}=1-K^{1-n}.$ (17)
Given assumption (A1), whenever two distinct points are selected, their distance is at least $\sigma_{\min}$. Therefore:
$\mathbb{E}_{\hat{q}}[D(\{x^{\prime}_{i}\}_{i=1}^{n})]\geq\sigma_{\min}\cdot P_{n}(K)=\sigma_{\min}(1-K^{1-n}).$ (18)
Step 2: Relating KL divergence to total variation distance. By Pinsker’s inequality [csiszar1967information, tsybakov2009introduction], which states that $\|p-q\|_{\text{TV}}\leq\sqrt{\frac{1}{2}D_{\text{KL}}(q\|p)}$ for any two probability distributions $p$ and $q$, the total variation distance is bounded by the KL divergence:
$\|p_{\theta}(\cdot|x,y^{\prime},p,m)-\hat{q}(\cdot|x,y^{\prime},d_{p,m})\|_{\text{TV}}\leq\sqrt{\frac{1}{2}D_{\text{KL}}(\hat{q}\|p_{\theta})}\leq\sqrt{\frac{\epsilon}{2}}.$ (19)
Step 3: Bounding the difference in expected diversity. The diversity measure $D(\{x^{\prime}_{i}\}_{i=1}^{n})$ is bounded above by $\Delta$ (assumption A3). Applying Lemma A.1:
$|\mathbb{E}_{p_{\theta}}[D]-\mathbb{E}_{\hat{q}}[D]|\leq 2\cdot\Delta\cdot\|p_{\theta}-\hat{q}\|_{\text{TV}}\leq 2\Delta\sqrt{\frac{\epsilon}{2}}=\Delta\sqrt{2\epsilon}.$ (20)
Step 4: Deriving the lower bound. Considering the worst-case scenario where $p_{\theta}$ produces less diversity than $\hat{q}$:
$\mathbb{E}_{p_{\theta}}[D(\{x^{\prime}_{i}\}_{i=1}^{n})]\geq\mathbb{E}_{\hat{q}}[D(\{x^{\prime}_{i}\}_{i=1}^{n})]-\Delta\sqrt{2\epsilon}.$ (21)
Substituting the result from Step 1 yields the desired bound:
$\mathbb{E}_{p_{\theta}}[D(\{x^{\prime}_{i}\}_{i=1}^{n})]\geq\sigma_{\min}(1-K^{1-n})-\Delta\sqrt{2\epsilon}.$ (22)
A.3 Implications
Diversity Guarantee
Theorem A.2 establishes that DiCoFlex maintains strong diversity guarantees. For sufficiently large $K$, the probability term $(1-K^{1-n})$ approaches 1, ensuring that the expected diversity is approximately $\sigma_{\min}-\Delta\sqrt{2\epsilon}$. The theorem shows that the learned distribution $p_{\theta}$ recovers nearly the full empirical diversity, with deviation bounded by the training error $\epsilon$.
Impact of Training Quality
As training progresses and $\epsilon\rightarrow 0$, the perturbation term $\Delta\sqrt{2\epsilon}$ vanishes, and the diversity of counterfactuals generated by the learned flow converges to that of the empirical distribution. This theoretical result validates the practical observation in Section 4 that DiCoFlex achieves superior diversity metrics compared to baseline methods.
Sample Efficiency
The bound demonstrates that generating $n$ samples from DiCoFlex provides diversity guarantees that scale favorably with the number of neighbors $K$ used during training. For fixed $n$ and increasing $K$, the probability of obtaining diverse samples $(1-K^{1-n})$ approaches 1 exponentially fast, explaining the effectiveness of our KNN-based training approach.
A.4 Validity Guarantees
Our method provides inherent validity guarantees by construction. Since $\hat{q}(x^{\prime}|x,y^{\prime},d)$ (Equation (6)) exclusively samples from the set of $K$-nearest neighbors $N(x,y^{\prime},d,K)$, where $h(x_{i})=y^{\prime}$, we have the following deterministic property:
Proposition A.3 (Validity).
For all training samples drawn from $\hat{q}(x^{\prime}|x,y^{\prime},d)$, the classification model returns the target class with probability one:
$p(y^{\prime}|x^{\prime})=1,\quad\forall x^{\prime}\sim\hat{q}(x^{\prime}|x,y^{\prime},d),$ (23)
where $p(y^{\prime}|x^{\prime})$ is the class probability returned by the classification model $h(\cdot)$.
Proof.
By definition of $\hat{q}$ in Equation (6), we have $\hat{q}(x^{\prime}|x,y^{\prime},d)>0$ only if $x^{\prime}\in N(x,y^{\prime},d,K)$. The neighborhood $N(x,y^{\prime},d,K)$ consists exclusively of training examples for which $h(x^{\prime})=y^{\prime}$ by construction. Therefore, $p(y^{\prime}|x^{\prime})=1$ for all samples from $\hat{q}$. ∎
This theoretical property explains why DiCoFlex achieves high validity in all datasets in Table 1 while being orders of magnitude faster than optimization-based methods. The normalizing flow trained on these valid samples inherits this property in expectation:
$\mathbb{E}_{x^{\prime}\sim p_{\theta}}[p(y^{\prime}|x^{\prime})]\approx 1,$ (24)
with the approximation quality depending on the training convergence (assumption A2 in Theorem A.2).
A.5 Proximity Guarantees
Our KNN sampling mechanism provides an implicit theoretical guarantee on proximity without requiring explicit proximity optimization.
Proposition A.4 (Proximity Upper Bound).
For any counterfactual $x^{\prime}$ sampled from $\hat{q}(x^{\prime}|x,y^{\prime},d)$ (Equation (6)), we have:
$d(x,x^{\prime})\leq d(x,x_{K}^{y^{\prime}}),$ (25)
where $x_{K}^{y^{\prime}}$ is the $K$-th nearest neighbor of $x$ in class $y^{\prime}$ under distance measure $d(\cdot,\cdot)$.
Proof.
By definition of $\hat{q}$ in Equation (6), we have $\hat{q}(x^{\prime}|x,y^{\prime},d)>0$ if and only if $x^{\prime}\in N(x,y^{\prime},d,K)$. The set $N(x,y^{\prime},d,K)$ contains exactly the $K$ nearest neighbors of $x$ in class $y^{\prime}$ according to distance $d$. Therefore, any $x^{\prime}$ with non-zero probability under $\hat{q}$ must satisfy:
$d(x,x^{\prime})\leq\max_{x_{i}\in N(x,y^{\prime},d,K)}d(x,x_{i})=d(x,x_{K}^{y^{\prime}}),$ (26)
as otherwise $x^{\prime}$ would not be among the $K$ nearest neighbors and would have $\hat{q}(x^{\prime}|x,y^{\prime},d)=0$. ∎
This provides an upper bound on proximity without requiring iterative optimization. The learned flow $p_{\theta}$ approximately respects this bound, with deviations controlled by the training error $\epsilon$ (assumption A2). The flexibility to adjust the distance measure $d_{p,m}$ (Equation (8)) allows users to control proximity characteristics at inference time through the sparsity parameter $p$ and mask $m$.
A.6 Plausibility Guarantees
We now establish theoretical guarantees for plausibility, addressing a key limitation of optimization-based methods that may generate out-of-distribution counterfactuals.
Proposition A.5 (In-Distribution Guarantee).
Since the training distribution $\hat{q}(x^{\prime}|x,y^{\prime},d)$ only assigns non-zero probability to $K$-nearest neighbors from the training set $\mathcal{D}$ (Equation (6)), any sampled counterfactual $x^{\prime}$ satisfies:
$p_{\text{data}}(x^{\prime})\geq\min_{x_{i}\in\mathcal{D}}p_{\text{data}}(x_{i})>0,$ (27)
where $p_{\text{data}}$ denotes the true data distribution.
Proof.
By construction, $\hat{q}(x^{\prime}|x,y^{\prime},d)$ assigns uniform probability $\frac{1}{K}$ to each of the $K$ neighbors in $N(x,y^{\prime},d,K)\subset\mathcal{D}$ and zero probability elsewhere. Therefore, any sample $x^{\prime}\sim\hat{q}$ is an actual training point: $x^{\prime}\in\mathcal{D}$. Since all training points are drawn from $p_{\text{data}}$, we have $p_{\text{data}}(x^{\prime})>0$ for all such $x^{\prime}$. ∎
Enhanced Plausibility with Sparse Constraints
When using small $p$ values (e.g., $p=0.01$) in the $L_{p}$ norm distance (Equation (7)), we restrict the neighborhood to samples with minimal feature changes. These neighbors typically lie in higher-density regions of the data manifold, as points with few modified features are more likely to remain within dense areas of the feature space. This explains the improved Local Outlier Factor (LOF) scores for DiCoFlex (p=0.01) compared to DiCoFlex (p=2.0) observed in Table 1.
The normalizing flow learns to interpolate between these guaranteed in-distribution points while preserving the manifold structure through its invertible transformations. This provides a principled approach to ensuring plausibility, as the flow is constrained to learn smooth deformations between verified in-distribution samples rather than arbitrary transformations that might venture into low-density regions.
Appendix B Dataset Descriptions